How many pound of hamburger that cost $1.60 per pound must be mixed with 70 pounds of hamburger that cost $2.10 per pound to make a mixture that cost $1.70 per pound.
Answer by ptaylor(1894)

Let x=amount of $1.60 hamburger needed
($ understood)
Value of the 1.60 hamburger=1.60x
Value of the 2.10 per pound hamburger=2.10*70
Now we know that the value of the 1.60 per pound hamburger(1.60x) plus the value of the 2.10 per pound hamburger (2.10*70) has to equal the value of the final mixture (1.70(70+x)), so our equation to solve is:
1.60x +2.10*70=1.70(70+x) get rid of parens (distributive law) and simplify
1.60x+147=119+1.70x subtract 1.60x and also 119 from each side
1.60x-1.60x+147-119=119-119+1.70x-1.60x collect like terms
28=0.10x divide each side by 0.10
x=280 lb-----------------------------amount of 1.60 per lb hamburger needed
CK
1.60*280+147=1.70*350
448+147=595
595=595
Hope this helps---ptaylor